### v3

# Evaluating the Defense Potential of Machine Unlearning against Membership Inference Attacks

Abstract
Membership Inference Attacks (MIAs) pose a significant privacy risk, as they enable adversaries to determine whether a specific data point was included in the training dataset of a model. While Machine Unlearning is primarily designed as a privacy mechanism to efficiently remove private data from a machine learning model without the need for full retraining, its impact on the susceptibility of models to MIA remains an open question. In this study, we systematically assess the vulnerability of models to MIA after applying state-of-art Machine Unlearning algorithms. Our analysis spans four diverse datasets (two from the image domain and two in tabular format), exploring how different unlearning approaches influence the exposure of models to membership inference. The findings highlight that while Machine Unlearning is not inherently a countermeasure against MIA, the unlearning algorithm and data characteristics can significantly affect a model’s vulnerability. This work provides essential insights into the interplay between Machine Unlearning and MIAs, offering guidance for the design of privacy-preserving machine learning systems.
1 Introduction
As machine learning (ML) models continue to advance, their applications have become widespread across various industries, such as healthcare nikolaidis2023federated , telecommunications tsiolakis2024carbon , biological sciences pavlidis2023extensive and beyond. These models have revolutionized how we approach tasks, making decision-making processes more efficient and accurate. However, the extensive use of ML raises concerns about the reliance on large volumes of user data, often containing sensitive or personal information. With the increasing number of privacy incidents where unauthorized parties gain access to this sensitive data, the need for robust privacy-preserving mechanisms has become critical.
One of the primary privacy vulnerabilities of ML models is defending against Membership Inference Attacks (MIAs) shokri2017membership . In such attacks, adversaries aim to determine whether a specific data point was part of a model’s training dataset. This information can potentially lead to the leakage of sensitive data, posing a significant threat to user privacy, especially when sensitive data such as medical records are involved. For instance, if a healthcare model can be queried to confirm whether specific patient records were used during its training, it could expose confidential health information about individuals. The implications of such attacks extend to other domains such as social media, finance and security. Therefore, understanding and mitigating the risks posed by MIAs is crucial for ensuring the privacy and safety of users.
In recent years, Machine Unlearning Cao has gained attention as a potential method for addressing privacy concerns in ML applications. Machine Unlearning allows service providers to erase specific data from a trained model without needing to retrain the model from scratch. This technique is particularly useful in the context of privacy regulations like the General Data Protection Regulation (GDPR), which requires the deletion of user data upon request. By using such erasing algorithms, organizations can efficiently comply with such legal requirements while minimizing computational costs and time associated with full model retraining. Furthermore, Machine Unlearning can be utilized as a security measure zhang2023review , since it reduces the influence of deleted data points on model predictions, potentially lowering the model’s susceptibility to adversarial attacks like MIAs.
Despite its appeal, the relationship between Machine Unlearning and a model’s vulnerability to MIAs remains largely unexplored. While Machine Unlearning primarily focuses on the efficient removal of data to protect user privacy, it is unclear how this process influences the model’s susceptibility to MIAs liu2024threats . This knowledge gap is crucial because even after certain data has been ”unlearned”, models may still retain enough information about the deleted samples to be vulnerable to MIAs. As such, merely applying unlearning techniques may not suffice to mitigate these risks unless the process is carefully evaluated and tailored to address the specific threats posed by MIAs.
This work aims to fill this gap by assessing the impact of Machine Unlearning on a model’s vulnerability against MIAs. We evaluate the efficacy of state-of-the-art Machine Unlearning algorithms across four distinct datasets, each representing different applications. Our analysis focuses on how unlearning methods affect a model’s exposure to MIAs and whether certain unlearning strategies are more effective than others in reducing this risk. Our findings suggest several important insights. First, while Machine Unlearning can effectively remove targeted data, it does not inherently serve as a defense mechanism against MIAs. In some cases, the removal of data via unlearning may even introduce subtle vulnerabilities that could increase the model’s susceptibility to adversarial attacks. Second, the effectiveness of unlearning techniques in reducing vulnerability to MIAs varies significantly depending on the unlearning algorithm used and the use-case scenario. Hence, designing a robust Machine Unlearning solution requires careful consideration of the the specific nature of the data being unlearned.
The rest of this paper is structured as follows: Section 2 defines key concepts, including MIAs and Machine Unlearning, and explores the influence of unlearning algorithms on MIA vulnerabilities. Section 3 presents related work, focusing on MIA, machine unlearning techniques, and the interplay between the two. In Section 4, we describe the datasets used in this study, covering both image and tabular data. Section 5 details the experimental setup, discussing the models and unlearning algorithms. Section 6 we present the results across different datasets, along with a sensitivity analysis. Finally, Section 7 concludes the paper with a summary of the findings and suggestions for future work.
2 Problem Definition
2.1 Membership Inference Attacks (MIAs)
MIAs present a significant privacy challenge in ML models. Specifically, a MIA involves an adversary trying to determine if a specific data point was part of the model’s training dataset hu2021membership . Given access to the model’s outputs, the adversary can exploit differences in the model’s behavior when given data that was part of the training set versus data that was not (see Fig. 1). This capability enables the adversary to deduce whether specific instances are in the training set, potentially revealing sensitive information about individuals.
Figure 1: MIA - Flow Diagram
Formally, let $\mathcal{D}_{train}$ represent the training dataset used to train a model $f$, and $\mathcal{D}_{test}$ represent a separate test dataset. The goal of an adversary is to design a membership inference attack algorithm, $A$, that, given a data point $x$ and model $f$, can output a binary label:
$A(f(x))=\begin{cases}1&\text{if }x\in\mathcal{D}_{train}\\ 0&\text{if }x\notin\mathcal{D}_{train}\end{cases}$
The success of a MIA depends on how effectively the adversary can differentiate between members of $\mathcal{D}_{train}$ and non-members based on the model’s predictions. Such attacks pose significant privacy risks as it could potentially leak sensitive information, especially in models trained on personal or confidential data, such as medical or financial records. In this work, we implement a white-box attack, where the adversary has knowledge of the target model’s architecture and the distribution of the dataset niu2024survey . This type of attack is more powerful, as it exploits deeper insights into the model, making it easier to infer membership.
2.2 Machine Unlearning
Machine Unlearning refers to the efficient process of removing the influence of specific data points from a trained ML model. This concept is particularly important in situations where users request the deletion of their data, especially under privacy regulations like the General Data Protection Regulation (GDPR) 10488864 . The challenge is to ensure that, after the data is removed, the model no longer retains any trace of the deleted data, while avoiding the expensive process of retraining the entire model from scratch (see Fig. 2).
Figure 2: Machine Unlearning - Flow Diagram
Let $\mathcal{D}_{train}=\mathcal{D}_{retain}\cup\mathcal{D}_{forget}$, where $\mathcal{D}_{forget}$ represents the data that needs to be unlearned. Given a model $f$ trained on $\mathcal{D}_{train}$, the goal of Machine Unlearning is to produce a modified model $f^{\prime}$ such that:
$f^{\prime}(x)\approx f(x)\quad\forall x\in\mathcal{D}_{retain}$
and the influence of $\mathcal{D}_{forget}$ is effectively removed. In other words, for any data point $x\in\mathcal{D}_{forget}$, the model $f^{\prime}$ should behave as though $x$ was never part of the original training dataset. Achieving this efficiently, without retraining the model from scratch, is the core challenge of Machine Unlearning.
2.3 Influence of Machine Unlearning on Membership Inference Attacks (MIAs)
While Machine Unlearning is designed to remove data points from a model’s training set, its influence on the model’s vulnerability to MIA is not straightforward. The process of unlearning aims to erase the impact of the data to be forgotten ($\mathcal{D}_{forget}$), but this can either reduce or exacerbate the model’s susceptibility to MIAs depending on how the unlearning is implemented Chen .
The key question is whether the model $f^{\prime}$ after unlearning becomes less distinguishable with respect to the forget data points, thus mitigating the effectiveness of MIAs, or whether the unlearning process itself leaves traces that adversaries can exploit. In practice, if the model retains subtle information about $\mathcal{D}_{forget}$ even after unlearning, it might still be vulnerable to MIAs, as adversaries could use this residual information to determine whether specific data points were in the original training set.
Formally, we are concerned with the impact of Machine Unlearning on the adversary’s success rate, $S(A,f^{\prime})$, where:
$S(A,f^{\prime})=$ $\,P(A(f^{\prime}(x))=1\mid x\in\mathcal{D}_{forget})$
$+$ $\,P(A(f^{\prime}(x))=0\mid x\notin\mathcal{D}_{forget}),$ (1)
where $P(A(f^{\prime}(x))=1)$ denotes the probability that the adversary’s MIA algorithm $A$ predicts that the data point $x$ was part of the original training dataset $\mathcal{D}_{train}$, based on the modified model $f^{\prime}$ after the unlearning process. Specifically, the adversary queries the model $f^{\prime}$ with $x$ and, based on the model’s output and attempts to infer whether $x$ was included in the forget dataset $\mathcal{D}_{forget}$. A prediction of $1$ indicates that the adversary believes $x$ was in the original training set.
Similarly, $P(A(f^{\prime}(x))=0)$ represents the probability that the adversary predicts that $x$ was not part of the original training set, indicating that $x$ was not in $\mathcal{D}_{forget}$. Together, these probabilities reflect the adversary’s success in conducting a membership inference attack on the unlearned model $f^{\prime}$.
The goal is for $S(A,f^{\prime})$ to be no better than random guessing for adversaries after the unlearning process, ideally reducing their success rate to that of flipping a coin (i.e., $S(A,f^{\prime})\approx 0.5$). However, depending on the unlearning algorithm, $S(A,f^{\prime})$ could remain higher, indicating that the model is still vulnerable to MIAs. Understanding and mitigating this influence is critical for designing robust unlearning mechanisms that not only comply with data deletion requests but also enhance the overall privacy and security of the model.
3 Related Work
3.1 MIA
Shokri et al. shokri2017membership introduced the concept of MIAs, demonstrating that adversaries could determine whether a particular data point was part of the training dataset by observing the model’s outputs. Their work laid the groundwork for subsequent research in this domain.
Salem et al. salem2018ml extended the MIA concept to more practical scenarios, showing that even with limited knowledge of the target model, adversaries could perform effective MIA. They explored different attack models, including black-box and white-box scenarios, highlighting the pervasive risk of MIA.
A comprehensive survey by Hu et al. hu2021membership details the variety of MIA techniques across different ML models, including classification and generative models. Their work also outlines the defences proposed to mitigate these attacks, emphasizing the ongoing need for robust privacy-preserving mechanisms. Other studies, such as Zhong2022Understanding explore the privacy unfairness in MIA and how certain defence mechanisms, like differential privacy and dropout, affect different subgroup data.
Further advancements in MIA, such as the MIA-BAD framework proposed by Banerjee et al banerjee2023miabad leverage federated learning as a potential defence mechanism against batch-wise MIA attacks. Similarly, Shi et al. shi2023scalemia introduce Scale-MIA, a scalable attack capable of recovering training data even in secure federated learning environments, demonstrating the persistent risks to user privacy.
3.2 Machine Unlearning
Cao et al. Cao introduced the foundational concept of machine unlearning, where they focused on the efficient removal of data from ML models. Their method transforms models like Naive Bayes into a summation-based structure, enabling the deletion of specific data points without requiring full retraining of the entire model.
Building on this foundational work, Bourtoule et al. Bourtoule introduced a more scalable unlearning framework called SISA in 2019. The core idea behind SISA training is to divide the training dataset into several disjoint shards and train sub-models on these separate shards. This approach enhances scalability since, when a data point needs to be removed, only the sub-models trained on the affected shards need retraining.
Ginart et al. Ginart tackled unlearning in the context of unsupervised learning, specifically k-means clustering. Their method quantizes data points into discrete clusters, enabling efficient unlearning by allowing modifications or deletions of the clusters containing the data points in question, without needing to retrain the entire model.
While these approaches focus on unlearning efficiency, it is equally important to minimize the negative impact on model accuracy. Addressing this concern, Perifanis et al. perifanis2024sftc proposed SFTC, an approach that fine-tunes the model on the remaining instances while intentionally confusing the model about the forget set, thus balancing unlearning with preserving predictive performance.
Moving to the context of logit-based classifiers, Baumhauer et al. Baumhauer explored machine unlearning by leveraging the structure of logit-based models. This approach provides an efficient method for removing data points, avoiding full retraining while ensuring data removal.
To tie all these efforts together, Xu et al. Xu conducted a comprehensive survey of various machine unlearning techniques. Their survey covers the broad range of methods introduced so far, analyzing the trade-offs between accuracy, efficiency, and scalability, and highlighting the ongoing challenges in this evolving field.
3.3 Machine Unlearning and MIA
Machine unlearning has emerged as a promising defense mechanism against MIAs, where adversaries attempt to infer whether specific data points were part of the training set. Several studies have explored various aspects of unlearning in this context, shedding light on both its potential and its limitations.
For instance, Sommer et al. Sommer investigate probabilistic verification methods to defend against MIA in unlearning systems, emphasizing the need for strong verification protocols to prevent data leakage. Their work underscores the necessity of robust mechanisms to ensure the integrity of unlearning processes.
Also, Doughan and Itani DoughanItani provide a comparative analysis of various machine unlearning methods, focusing on their effectiveness in hindering MIAs. They emphasize the importance of future research in this area, particularly in the context of different models and datasets. In their study, they examine machine unlearning techniques applied to Resnet-18 models trained with the CIFAR-10 dataset. By comparing Fine-Tuning and Fisher Noise-based Impair-Repair methods, they demonstrate that the latter significantly reduces MIA scores while maintaining model accuracy.
However, even with such advancements, new challenges continue to emerge. Lu et al. Lu introduce label-only MIAs, highlighting a novel threat to machine unlearning systems. These attacks enable adversaries to infer membership without relying on posterior probabilities, thus complicating existing defense strategies.
In addition to these concerns, Chen et al. Chen delve deeper into the privacy risks associated with unlearning. They show that MIAs can still exploit unlearning processes, even when sensitive data is expected to be removed. The authors propose a novel MIA that targets models post-unlearning by leveraging differences in model outputs before and after the unlearning process, demonstrating how attackers can infer which data points were likely part of the original training set.
Finally, Niu et al. NIU2024404 offer a comprehensive review of the landscape of MIAs. Surveying over 200 publications, they systematically examine various defense strategies, including machine unlearning, to mitigate these attacks. Their findings highlight the importance of careful evaluation in the unlearning process, noting that incomplete or ineffective unlearning can leave models vulnerable to residual traces of removed data.
Although the recent advances in the connection of unlearning to MIAs have expanded our understanding of defense mechanisms, most efforts focus mainly on image classification tasks, with limited attention given to other types of data, such as tabular datasets. In addition, most recent approaches tend to evaluate the effectiveness of MIAs by comparing only two snapshots of the model: one before unlearning and one after. While this evaluation provides a basic measure of improvement or degradation in MIA resistance, it neglects to offer insights into the dynamics of the unlearning process itself and how it incrementally impacts the model’s vulnerability to MIAs. To address these gaps, we conduct a comprehensive experimental study using both image and tabular datasets, evaluating MIAs not only before and after unlearning but throughout the entire unlearning process.
4 Datasets
To evaluate the defense potential of machine unlearning against MIA, we conduct experiments on four widely used datasets: CIFAR-10, Machine Unlearning Faces (MuFac), Purchase, and Texas100. The first two datasets are from the image domain, with Purchase and Texas100 be in tabular format. The four datasets represent sensitive real-world information, allowing us to assess the generalization and effectiveness of unlearning techniques in mitigating MIA across different domains. Below, we provide a detailed description of each dataset and its relevance to our experiments on machine unlearning.
4.1 Image Data
The CIFAR-10 dataset cifar10 is a popular benchmark dataset used in ML for image classification tasks. It consists of 60,000 color images, each of 32x32 pixels, categorized into 10 different classes. These classes are: airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks. Each class has 6,000 images, with 50,000 images designated for training and 10,000 for testing. The dataset is widely used due to its diversity and relatively small image size, making it a suitable challenge for various image recognition models, particularly in deep learning.
The MuFac Dataset choi2023towards is a multi-class age classification dataset, featuring over 13,000 Asian facial images with annotations for age groups and personal identities, ideal for machine unlearning research. The dataset provides a comprehensive collection of facial images with multiple factors influencing the appearance, including different angles, accessories (like glasses or hats), and emotions.
4.2 Tabular Data
The Purchase-100 dataset purchase100 is a customer purchase dataset widely used in privacy and adversarial ML research. It consists of 100 product classes, with each class representing a different product category. The dataset contains 600 binary features for each customer, indicating whether the customer purchased a specific product within a category. It is often used to study MIAs, making it useful for evaluating the privacy of ML models.
The Texas-100 dataset shokri2017membership , is a hospital discharge dataset that contains data from Texas hospitals. It includes 100 classes representing different diagnosis categories, with over 6,000 features per record. The features include patient demographics, treatment details and other hospital-related data. The Texas-100 dataset is commonly used in privacy research because of the sensitive nature of the information it contains, such as patient discharge status and procedures performed.
5 Experiments
In this section, we present the experimental setup and results of our study on the defense potential of machine unlearning algorithms against MIAs. Our experiments are designed to assess the unlearning performance of various state-of-the-art algorithms and evaluate their impact on different ML model architectures. We explore how effectively unlearning algorithms remove the influence of specific data points from trained models and how this influences the vulnerability of the models to MIAs. We also evaluate the utility of the models post-unlearning, ensuring that the removal of data does not significantly degrade predictive performance.
5.1 ML Model
All models were trained using an 80-20 train-test split. For each dataset and to evaluate the trained models against MIA, the target and shadow models used identical configurations, using five shadow models. Both target and shadow models used Multilayer Perceptron (MLP) architectures, for all datasets.
The training data for both the target and shadow models was divided equally, with a 50-50 split between target and shadow training data. MIAs were conducted using attack models that had equal proportions of members and non-members. The unlearning experiments followed the same data splitting as the training process, with an additional split to differentiate the retain and forget set (80-20 retain-forget split).
Table 1 summarizes the architectures used for each dataset, with MLP models employed across all datasets.
Each MLP architecture was specifically designed to balance efficiency and performance. The two fully connected (FC) layers with 512 and 256 units ensured that the models could handle the complexity of the image and tabular data while maintaining computational efficiency. The unlearning process, whether on specific classes or randomly selected rows, was designed to minimize the impact on retained data, ensuring that the models continued to perform well after unlearning.
Dataset Epochs Learning Rate Layers and Units CIFAR-10 30 0.001 3 FC layers (512, 256, 128 units) MuFac 60 0.0007 3 FC layers (512, 256, 128 units) Purchase-100 100 0.01 1 FC layers (128 units) Texas-100 100 0.01 1 FC layers (128 units)
Table 1: ML Model Architectures
5.2 Unlearning Algorithms
In this paper, we used several machine unlearning methods, and we compared their capacity to address MIAs. Our goal is to achieve a better understanding of these algorithms, while trying to understand the potential gaps and strengths of unlearning algorithms against this type of attacks.
We include the SFTC algorithm perifanis2024sftc as our main focus, given its novel approach to balancing utility preservation and targeted forgetting. In addition we use NegGrad, the most straightforward baseline which modifies model parameters by applying gradient ascent on the forget set, aiming to erase the influence of the forget set. We also consider SCRUB, a teacher-student approach, which minimizes the divergence between the student and original models on the retain set while maximizing divergence on the forget set. Each algorithm is tested under identical conditions to ensure a fair comparison of their unlearning and defense capabilities.
6 Results
6.1 CIFAR-10 Dataset
6.1.1 NegGrad Algorithm
We begin by examining the NegGrad algorithm applied to the CIFAR-10 dataset, which yields several insights into the unlearning process. A key highlight here is that when the forget accuracy reaches zero, it signifies that the model has successfully forgotten the specified data, as illustrated in Fig 3. Additionally, during the unlearning process, we observe a decline in test accuracies. This highlights a trade-off inherent in the algorithm: while the unlearning process effectively achieves perfect forgetting, it comes at the cost of reduced overall model performance.
Figure 3: ML Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm NegGrad
Furthermore, the Fig 4 illustrates the accuracy scores of the MIA attack on the forget and retained data, as the model undertakes the unlearning process per epoch. The accuracy for the forget data (solid blue line) shows a consistent decline, indicating that the unlearning process is effectively reducing the model’s ability to recall the forget data points, which aligns with the intended goal of making the model forget specific data. Conversely, the accuracy for the retained data (dashed orange line) also decreases over time, though at a slower pace initially. This outcome was expected, as the unlearning algorithm not only reduces the model’s ability to recall forget data but also impacts overall model performance, which in turn lowers the effectiveness of MIA attacks on both forget and retained data.
Figure 4: MIA Accuracy on CIFAR-10 Dataset / Unlearning Algorithm NegGrad
6.1.2 SCRUB Algorithm
On the other hand, the scrub algorithm applied to the CIFAR-10 dataset did not demonstrate the same functionality as the NegGrad approach. As depicted in Fig 5, the forget accuracy remained stable, showing no significant decline across epochs. Additionally, in the context of MIAs (see Fig. 6), we observed a slight decrease in the forget accuracy score. This suggests that the unlearning process may have some impact on mitigating the effectiveness of MIA attacks.
Figure 5: ML Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SCRUB
Figure 6: MIA Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SCRUB
6.1.3 SFTC Algorithm
In Figure 7 the results show that the SFTC algorithm on the CIFAR-10 dataset maintains high performance on the training data, with stable test and validation accuracies. However, the forget accuracy remains high, suggesting that the algorithm is not effectively unlearning the forget data. Compared to the other two algorithms, SFTC performs better in terms of maintaining model utility and generalization ability, making it the best option among the three for balancing unlearning and performance.
Figure 7: ML Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SFTC
Figure 8 shows the MIA accuracy trends for SFTC, where the forget accuracy steadily decreases while the retain accuracy remains relatively stable. This result suggests that SFTC effectively mitigates MIA risks by successfully unlearning forget data while maintaining model performance on retained data. Furthermore, this algorithm outperforms the others in mitigating MIA attacks, as evidenced by the lower MIA accuracy on the retained dataset compared to the other algorithms. Only the NegGrad algorithm achieved better (lower) MIA attack accuracies, but it does so at the cost of also degrading the performance on the retained dataset. This makes SFTC more effective overall, as it balances successful unlearning with maintaining model performance on the retained data.
Figure 8: MIA Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SFTC
6.2 MuFac Dataset
6.2.1 NegGrad Algorithm
The NegGrad algorithm performs significantly better on the MuFac dataset compered to CIFAR-10, as demonstrated by the Figure 9. This can be attributed to the simpler and more structured nature of the MuFac dataset, which likely consists of facial images with less variability between classes. Furthermore, the near-identical accuracy trends in both the training and test sets during the epochs for the MuFac dataset suggest that the model maintained stability. Lastly, the clear gap between the training accuracy, which remains at nearly 1.0, and the test accuracy, which hovers around 0.6, indicates significant overfitting.
Figure 9: ML Model Accuracy on MuFac Dataset / Unlearning Algorithm NegGrad
Figure 10 presents the MIA accuracy trends, where the forget accuracies decline sharply over the epochs, indicating that NegGrad is effectively unlearning the forget data with respect to MIA attacks. This declining trend in forget accuracies aligns with the goal of reducing the model’s ability to recall forget data. Meanwhile, the retain accuracies remain relatively stable throughout the unlearning process, suggesting that the model’s ability to retain and accurately classify non-forget data is preserved. This balance shows that NegGrad is successfully mitigating MIA risks for forget data without significantly compromising the accuracy of retained data.
Figure 10: MIA Model Accuracy on MuFac Dataset / Unlearning Algorithm NegGrad
6.2.2 SCRUB Algorithm
Figure 11 shows the SCRUB algorithm’s performance on MuFac, where the forget accuracy shows a minor decline but remains relatively stable throughout the epochs. The validation and test accuracies also remain stable, indicating that the algorithm preserves model performance but may not be as effective in unlearning forget data.
Figure 11: ML Model Accuracy on MuFac Dataset / Unlearning Algorithm SCRUB
In Figure 12, the MIA accuracy trends for the SCRUB algorithm on the MuFac dataset are shown. The forget accuracies fluctuate considerably across the epochs but remain relatively high, hovering around 0.62-0.64. This suggests that the SCRUB algorithm is less effective at consistently unlearning the forget data, as the model still maintains a relatively high ability to recall forget information. Meanwhile, the retain accuracies also show fluctuations but remain slightly lower than the forget accuracies, mostly around 0.60-0.62. This indicates that the model’s ability to retain data is not significantly stronger than its ability to recall forget data, which suggests a trade-off between unlearning forget data and maintaining the retained data.
Figure 12: MIA Model Accuracy on MuFac Dataset / Unlearning Algorithm SCRUB
6.2.3 SFTC Algorithm
In Figure 13 the SFTC algorithm applied to the MuFac dataset shows high training accuracy, consistently close to 1.0, indicating that the model fits the training data extremely well. However, the test accuracy fluctuates around 0.3, revealing a significant gap between the training and test performance, which indicates substantial overfitting. Furthermore, the forget accuracy hovers around 0.85, suggesting that the algorithm has not been fully effective in making the model forget the specified data, as it retains a high level of accuracy on the forget data.
Figure 13: ML Model Accuracy on MuFac Dataset / Unlearning Algorithm SFTC
In Figure 14 we observe the MIA attack accuracies for the SFTC algorithm on the MuFac dataset as the model unlearns over time. The forget accuracies remain relatively stable, fluctuating between 0.60 and 0.65 across the epochs. This indicates that the model’s vulnerability to MIA attacks on forget data is reducing, which is aligned with the unlearning objective, but the algorithm is not fully effective in completely mitigating MIA attacks on forget data. In contrast, the retain accuracies are higher, fluctuating between 0.70 and 0.75, suggesting that the model remains more vulnerable to MIA attacks on retained data. This higher and more variable accuracy for retained data suggests that while the SFTC algorithm successfully diminishes the model’s ability to recall forget data (as seen by the lower MIA accuracy), it still struggles to maintain a consistent level of resistance to MIA attacks on retained data.
Figure 14: MIA Model Accuracy on MuFac Dataset / Unlearning Algorithm SFTC
6.3 PURCHASE Dataset
6.3.1 NegGrad Algorithm
In Figure 15, we observe the performance of the NegGrad algorithm on the Purchase dataset, showing trends that closely mirror those observed with the CIFAR-10 dataset. The training accuracy and forget accuracy both start high, around 0.8, but decline steadily over the epochs, reaching approximately 0.1 by the end. This sharp decline indicates that the algorithm is effectively unlearning the specified data, as the model gradually loses its ability to recall the forget information, similar to its behavior on CIFAR-10. Furthermore, the results on the Purchase dataset confirm the pattern observed with CIFAR-10: the NegGrad algorithm effectively forgets targeted data, but this unlearning process leads to a notable decline in test and validation accuracies, indicating compromised generalization. This highlights the inherent trade-off in the NegGrad algorithm, where successful unlearning is achieved at the expense of the model’s ability to retain its performance on retained and unseen data.
Figure 15: ML Model Accuracy on Purchase Dataset / Unlearning Algorithm NegGrad
In Figure 16 the MIA attack accuracies for the NegGrad algorithm on the Purchase dataset demonstrate how the model’s vulnerability to MIA attacks changes during the unlearning process. The forget accuracies steadily decrease from approximately 0.58 to 0.48 over the course of 50 epochs, indicating the algorithm’s success in making the model forget specific data points. This declining trend highlights NegGrad’s effectiveness in reducing the model’s ability to recall forget data, thereby achieving the intended unlearning objective. However, as seen with other datasets, this process comes with a trade-off. The retain accuracies also decrease, though less sharply, from around 0.55 to 0.47, indicating that the model’s resistance to MIA attacks on retained data diminishes over time. This convergence between forget and retain accuracies suggests that, by the end of the unlearning process, the model’s vulnerability to MIA attacks becomes nearly the same for both forget and retained data. While NegGrad is effective in mitigating MIA risks for forget data, it also impacts the model’s overall performance and resistance to attacks on retained data, similar to its behavior on other datasets.
Figure 16: MIA Model Accuracy on Purchase Dataset / Unlearning Algorithm NegGrad
6.3.2 SCRUB Algorithm
Figure 17 demonstrate that the SCRUB algorithm is effective in unlearning forget data for the Purchase dataset, as evidenced by the sharp decline in forget accuracy. However, this unlearning comes at a cost: both the test and validation accuracies gradually decrease, suggesting that the model’s generalization ability on unseen data is compromised as the unlearning progresses. Unlike the results seen on the CIFAR-10 and MuFac datasets, the SCRUB algorithm performs better on the Purchase dataset, showing a more successful and consistent reduction in forget accuracies while maintaining high training accuracy. Nonetheless, the trade-off in generalization performance remains a consistent challenge across datasets.
Figure 17: ML Model Accuracy on Purchase Dataset / Unlearning Algorithm SCRUB
In Figure 18, show that the SCRUB algorithm performs well in reducing MIA accuracy for forget data, making it more difficult for adversaries to identify whether forget data was part of the model’s training set. However, the model’s susceptibility to MIA attacks on retained data is also affected, albeit to a lesser extent. Overall, the SCRUB algorithm effectively balances unlearning and retaining data while mitigating MIA risks, particularly for forget data.
Figure 18: MIA Model Accuracy on Purchase Dataset / Unlearning Algorithm SCRUB
6.3.3 SFTC Algorithm
Figure 19 indicate that the SFTC algorithm is ineffective in unlearning for the Purchase dataset. While the model maintains high training accuracy, the forget accuracy remains almost unchanged, indicating that the algorithm fails to reduce the model’s ability to recall the forget data. Additionally, the test and validation accuracies remain stable, implying that the overall generalization performance is not significantly affected by the unlearning process.
Figure 19: ML Model Accuracy on Purchase Dataset / Unlearning Algorithm SFTC
In Figure 20, shows that the SFTC algorithm effectively reduces MIA attack accuracy for forget data while maintaining relatively stable performance for retained data. This suggests that the algorithm strikes a good balance between unlearning and retaining data, successfully mitigating MIA risks for forget data without significantly compromising the model’s protection of retained data. The steady separation between forget and retain accuracies highlights the algorithm’s effectiveness in reducing the model’s ability to recall forget data while preserving its resilience for retained data.
An interesting observation here is the distinction between general unlearning performance and the model’s vulnerability to MIA attacks. In Fig 19, the high forget accuracy shows that the model hasn’t fully unlearned the forget data, meaning it can still predict it accurately. However, in Fig 20, the decline in MIA attack accuracy suggests that the model becomes more resistant to adversarial attacks, likely because the features it uses to classify the forget data are becoming less distinctive. This indicates that even when the unlearning process is incomplete, the model can still improve its resilience to MIA attacks.
Figure 20: MIA Model Accuracy on Purchase Dataset / Unlearning Algorithm SFTC
6.4 Texas100 Dataset
6.4.1 NegGrad Algorithm
In Figure 21 the NegGrad algorithm achieves near complete unlearning, as evidenced by the sharp drop in forget accuracy after 20 epochs. However, this comes at the cost of the model’s overall performance, as seen by the simultaneous collapse of training, test, and validation accuracies to 0. The results suggest that while NegGrad is very effective in unlearning, it severely impacts the model’s ability to retain useful information for both training and generalization purposes. This consistent behavior across datasets, except for the Purchase, highlights a key limitation of the NegGrad algorithm: while it is highly effective in achieving unlearning, it often does so at the expense of the model’s overall performance, leading to a complete collapse in its ability to retain useful information and generalize to new data.
Figure 21: ML Model Accuracy on Texas100 Dataset / Unlearning Algorithm NegGrad
Figure 22 shows that the NegGrad algorithm is highly effective in mitigating the model’s vulnerability to MIA attacks on both forget and retained data in the Texas dataset. After 30 epochs, the model reaches a state where it no longer provides useful information to an adversary trying to conduct an MIA attack, as both the forget and retain accuracies stabilize at around 0.5, equivalent to random guessing. This suggests that while the unlearning process is complete, the model’s ability to differentiate between forget and retained data has been eliminated, achieving strong resistance to MIA attacks across the board.
Figure 22: MIA Model Accuracy on Texas100 Dataset / Unlearning Algorithm NegGrad
6.4.2 SCRUB Algorithm
In Figure 23 the SCRUB algorithm demonstrates a controlled unlearning process, gradually reducing the model’s ability to recall forget data while maintaining stable generalization performance on the test and validation sets. The slow decline in forget accuracy shows that the unlearning is effective but not too aggressive, while the near-perfect training accuracy and stable test/validation accuracies suggest that the model continues to generalize well even as it forgets specific data. This indicates a well-balanced approach, where unlearning is achieved without drastically compromising the model’s overall performance.
Figure 23: ML Model Accuracy on Texas100 Dataset / Unlearning Algorithm SCRUB
In Figure 24 the SCRUB algorithm shows a successful reduction in MIA attack accuracy for forget data while maintaining relatively stable MIA attack accuracy for retained data. The steady decline in forget accuracies indicates that the model is becoming more resistant to MIA attacks on forget data, achieving effective unlearning. At the same time, the stability of retain accuracies suggests that the model’s protection for retained data remains largely intact. This balance between unlearning and retaining data highlights the SCRUB algorithm’s ability to effectively mitigate MIA risks for forget data without compromising the security of retained information.
Figure 24: MIA Model Accuracy on Texas100 Dataset / Unlearning Algorithm SCRUB
6.4.3 SFTC Algorithm
Figure 25 shows that the SFTC algorithm maintains high accuracy across all metrics, but this comes at the cost of its unlearning effectiveness. The forget accuracy remains high, indicating that the model retains its ability to recall the forget data, meaning the unlearning process is largely unsuccessful. However, the model’s test and validation accuracies remain stable, indicating that its generalization ability is preserved. Overall, the SFTC algorithm appears to preserve the model’s original performance but fails to achieve effective unlearning for the forget data in the Texas dataset.
Figure 25: ML Model Accuracy on Texas100 Dataset / Unlearning Algorithm SFTC
In Figure 26, the SFTC algorithm shows a successful reduction in MIA attack accuracy for forget data while keeping MIA accuracy for retained data stable. The steady decline in forget accuracies demonstrates that the model is becoming less vulnerable to MIA attacks on forget data as the unlearning process progresses. At the same time, the stability of retain accuracies suggests that the model continues to provide protection for retained data. This balance between unlearning and maintaining security for retained data highlights the SFTC algorithm’s effectiveness in managing MIA risks for forget data without compromising retained information.
Figure 26: MIA Model Accuracy on Texas100 Dataset / Unlearning Algorithm SFTC
6.5 Sensitivity Analysis
The sensitivity analysis of the model’s performance across various learning rates (see Fig 27) reveals a key trade-off between model accuracy stability and resilience to Membership Inference Attacks (MIA). The first chart demonstrates that, across all learning rates, the model achieves high accuracy for both forgotten and retained data, with accuracies consistently ranging from 0.97 to 1.0 for forgotten data and slightly lower but still stable accuracy for retained data. Lower learning rates, such as (0.0005), show minimal fluctuations in performance, indicating greater stability over time. Higher learning rates, while still producing high accuracy, introduce minor variability, suggesting that these may lead to faster learning at the cost of small performance fluctuations.
In contrast, the second chart (see Fig 28), focusing on MIA accuracy, highlights that higher learning rates significantly improve the model’s resistance to MIA attacks. The MIA accuracy for forgotten data declines more rapidly with higher learning rates, suggesting that these configurations reduce the model’s vulnerability to attacks more effectively. Conversely, lower learning rates, although stable in terms of accuracy, leave the model more susceptible to privacy risks, particularly for retained data where MIA accuracy remains higher. This analysis suggests that while lower learning rates offer more consistent accuracy, higher learning rates provide better protection against MIA attacks, especially for forgotten data, striking a crucial balance between stability and privacy.
Figure 27: sensitivity mia accuracy
Figure 28: sensitivity mia accuracy
7 Conclusion
This paper investigates the effectiveness of Machine Unlearning as a defense mechanism against MIAs across a range of datasets and model architectures. Our experimental results demonstrated that Machine Unlearning can indeed play a significant role in reducing a model’s vulnerability to MIAs, though the effectiveness of this defense is highly dependent on the choice of unlearning algorithm.
Among the algorithms evaluated, i.e., NegGrad, SCRUB and SFTC, the latter consistently outperformed the others in terms of both unlearning effectiveness and preservation of model utility. It achieved a successful balance by thoroughly erasing the influence of forget data while maintaining high accuracy on retained data. This aspect is particular important in privacy-sensitive applications, where compliance with regulations like the GDPR requires robust data removal, but without compromising the model’s predictive power.
NegGrad, although effective in unlearning the forget set, exhibited significant overfitting, leading to lower test accuracy and greater degradation in model performance. SCRUB, on the other hand, offered moderate success, preserving model utility but being less effective in fully forgetting the targeted data, as evidenced by relatively stable forget accuracy in several datasets. Both methods showed weaknesses when compared to SFTC, particularly in terms of maintaining generalization and reducing MIA risks.
The sensitivity analysis revealed that higher learning rates generally improved the effectiveness of unlearning, with SFTC proving less sensitive to learning rate changes than NegGrad or SCRUB. This robustness makes SFTC more adaptable and practical for real-world ML applications, where hyperparameter tuning may be limited.
Despite the promise of Machine Unlearning, it is clear that this technique is not a universal safeguard against MIAs. While SFTC shows considerable potential in mitigating the risks, the choice of unlearning algorithm and its interaction with specific datasets and architectures are critical factors in determining the overall level of protection. Moreover, residual vulnerabilities may still exist in models post-unlearning, as attackers could exploit subtle traces of the forget data.
While this study provides valuable insights into the defensive potential of machine unlearning against MIAs, there remain several areas for future exploration to further enhance its effectiveness. This includes investigating adaptive unlearning strategies, which dynamically adjust based on the model’s vulnerability to MIAs, and evaluating the impact of unlearning on other types of adversarial attacks. Additionally, integrating Machine Unlearning with other privacy-preserving mechanisms, such as differential privacy, could enhance the overall defense against MIAs.
It will also be important to explore the long-term effects of repeated unlearning requests and their impact on model stability, as well as to develop standardized benchmarks for evaluating the robustness of unlearning algorithms. As MIAs evolve, unlearning methods must continue to adapt, ensuring comprehensive protection for user data in an increasingly adversarial landscape.